# Interferometry of non-Abelian band singularities and Euler class topology

### v1

@@ Abstract
In systems with a real Bloch Hamiltonian band nodes can be characterised by a non-Abelian frame-rotation charge. The ability of these band nodes to annihilate pairwise is path dependent, since by braiding nodes in adjacent gaps the sign of their charges can be changed. Here, we theoretically construct and numerically confirm two concrete methods to experimentally probe these non-Abelian braiding processes and charges in ultracold atomic systems. We consider a coherent superposition of two bands that can be created by moving atoms through the band singularities at some angle in momentum space. Analyzing the dependency on the frame charges, we demonstrate an interferometry scheme passing through two band nodes, which reveals the relative frame charges and allows for measuring the multi-gap topological invariant. The second method relies on a single wavepacket probing two nodes sequentially, where the frame charges can be determined from the band populations. Our results present a feasible avenue for measuring non-Abelian charges of band nodes and the experimental verification of braiding procedures directly, which can be applied in a variety of settings including the recently discovered anomalous non-Abelian phases arising under periodic driving.
@@
Band nodes also play an important role in single-gap topological phases. Conventional Weyl nodes act as sources of Berry flux [30], while higher order degeneracies require stabilization of crystalline symmetries and are often associated with topological invariants. Individually, their properties can be probed in different experiments [31, 32, 33, 34, 35, 36]. For example, the $\pi$-Berry flux of a Dirac cone has been measured through atomic momentum-space interferometry in optical lattices [32], and the winding of isolated linear or quadratic band touchings has been observed via exciting atoms by moving them through the node [34]. Upon being elevated to multi-gap topologies, the hallmark gained by band singularities is non-Abelian frame charges that can induce a nontrivial Euler class. Consequently, a fundamental question arises whether there exists observable signatures of these non-Abelian charges, their braiding and this invariant.
@@
Non-Abelian Frame Charges and Euler Class.— For a real Hamiltonian, the eigenstates $|u_{n}(\bm{k})\rangle$ at quasimomentum $\bm{k}$ span an orthonormal triad or, in general, vielbein (’frame’). The space of spectralized Hamiltonians is given by $O(N)/O(1)^{N}=O(N)/\mathbb{Z}_{2}^{N}$ for $n\in N$-bands, where $N>2$. The quotient ensures that flipping the signs of the eigenvectors does not change the Hamiltonian. Band singularities obstruct a unique assignment of the eigenstates and correspond to a $\pi$-rotation between the eigenstates that host the node (see Fig. 1b). Namely, the frame rotation accumulates a $\pi$ Berry phase, where band nodes act as the analogues of $\pi$ disclination vortices in biaxial nematics [37, 38]. While within the two-band subspace (a single ‘gap’) the nodes act as $\pi$-rotations, where it should be noted that $2\pi$ rotations in fact correspond to $-1$ as $\pi_{1}(SO(N))=\mathbb{Z}_{2}$, accumulated frame charges anticommute with those in the adjacent gaps. Focusing on the case of three bands, the frame charges associated to the nodes $\pi_{1}(SO(3)/D_{2})=Q$ take values in the quaternion group $Q=\{\pm 1,\pm i,\pm j,\pm k\}$ [37]. As a result, braiding band nodes in momentum space (which can be achieved e.g. using stress/strain [17, 39], temperature effects [40] or periodic driving [29]), converts their charges and can ensure that a specific band subspace hosts similarly valued charges. In two dimensions, the resulting obstruction to annihilate these charges (nodes) between states $n$ and $(n+1)$ is quantified by the Euler class [17],
@@
evaluated over any patch $\mathcal{D}$ encapsulating these nodes. Here, we define the Euler form $\mathrm{Eu}=\langle\partial_{k_{1}}u_{n}(\boldsymbol{k})|\partial_{k_{2}}u_{n+1}(\boldsymbol{k})\rangle-\langle\partial_{k_{2}}u_{n}(\boldsymbol{k})|\partial_{k_{1}}u_{n+1}(\boldsymbol{k})\rangle$ and associated connection one-form, $\mathcal{A}=\langle u_{n}(\boldsymbol{k})|\boldsymbol{\nabla}u_{n+1}(\boldsymbol{k})\rangle$.
@@
The braiding and Euler class can effectively be captured by tracking Dirac strings [29]. These are gauge objects connecting pairs of nodes in each gap (see Fig. 1) and represent the line across which the sign of the eigenstates change due to the $\pi$-Berry phase induced by the band nodes. Crossing a Dirac string (DS) residing in an adjacent gap changes the sign of the frame charge, effectively encoding braiding rules of the band node charges.
@@
Model.— As a concrete example, we employ a $\mathcal{C}_{2}\mathcal{T}$-symmetric Kagome lattice, while our results indeed apply to any system admitting an Euler description. The Hamiltonian diagonal in momentum space is written as
@@
$H(\bm{k})=-2J\sum_{\beta\neq\beta^{\prime}}\cos{\left(\bm{k}\cdot\bm{d}_{\beta\beta^{\prime}}\right)}c^{\dagger}_{\beta}c_{\beta^{\prime}}+\sum_{\beta}\Delta_{\beta}c^{\dagger}_{\beta}c_{\beta},$ (2)
@@
with nearest-neighbour hopping of amplitude $J$ along the three directions $\bm{d}_{\beta\beta^{\prime}}$ connecting three sublattices $\beta\in(A,B,C)$, with the annihilation (creation) operators $c^{(\dagger)}_{\beta}$ [29]. For vanishing sublattice offsets $\Delta_{\beta}$, the spectrum features two linear band touching points which carry opposite frame charges (e.g. ±$\pm i$, depicted with empty/filled markers in Fig. 1b), hence a vanishing patch Euler class. The quadratic band touching point in gap 2 harbours $\chi=1$, corresponding to two same-valued frame charges which we can assign to be $+j$  [^1] The frame-charge assignment is a gauge choice, where only the relative sign per gap is physically meaningful.. While these quaternion frame charges act as ±$\pm\pi$ Berry fluxes in a given gap, conventional interferometry methods are insensitive to their signs and, hence, cannot a priori reveal the relative frame charges or their change after braiding.
Figure 1: a) Kagome lattice band structure for $\Delta_{\beta}=0$. b) Linear band nodes $(K,K^{\prime})$ carry opposite non-Abelian frame charges (±$(\pm i)$, and hence can annihilate ($\chi_{1,2}=0$ in the shaded area). They act as ±$\pm\pi$ fluxes in gap 1, captured by the DS connecting them. The $\Gamma$ node is formed by two similar-valued frame charges $(+j)$ that are obstructed to annihilate $(\chi_{2,3}=1)$. The orthonormal frame of the eigenstates (arrows) rotates around a band node, accumulating a $\pi$-phase jump at the DS. c) A wavepacket moving through a node at an angle $\theta$, and finishing in a coherent superposition [Eq. (Interferometry of non-Abelian band singularities and Euler class topology)] due to the frame rotation by an angle $\phi$.
@@
Method 1: Interferometry.— We employ a coherent superposition in the bands forming a topological singularity (i.e. the Euler subspace) to deduce the frame rotation around it relative to other nodes in this gap, hence, their relative frame charges. Consider moving a wavepacket in momentum space in band 1 $(\mathinner{|{u_{1,\mathrm{in}}(\bm{k})}\rangle})$, which enters and exits a band node at some angle $\theta$ as depicted in Fig. 1[34]. The frame of band eigenstates is rotated by an angle $\phi=\pm\theta/2$ between the in-going and out-going paths  [^2] For a uniform and symmetric Hamiltonian, the frame-rotation angle winds smoothly $\phi=\pm\theta/2\in[0,\pi]$ for the interferometry angle completing a circle $\theta\in[0,2\pi]$. Here $\pm$ indicates the direction of winding, hence the sign of the frame charge in the fixed gauge. In general, $\phi$ might change non-uniformly while still rotating by $\pi$ in total, see SM [44] for more elaboration.. The atoms, therefore, are in a coherent superposition after passing through the node (adiabatically with respect to other bands) as,
@@
$=\cos{\phi}\mathinner{|{u_{1,\mathrm{out}}}\rangle}-\sin{\phi}\mathinner{|{u_{2,\mathrm{out}}}\rangle},$ (3)
where the resulting probability to be in bands $\mathinner{|{u_{1,\mathrm{out}}}\rangle}$ and $\mathinner{|{u_{2,\mathrm{out}}}\rangle}$ is expressed in terms of the frame rotation $\phi$ corresponding to the overlap.
When the system is characterized by an Euler class within a fixed gauge, the relative phase of the component excited into the second band depends on the chirality of the frame winding, i.e. the full dreibein that is naturally encoded by the Hamiltonian evolution. By analyzing two such excitation events, we devise an interferometry of the Euler nodes to extract the relative non-Abelian frame charges and, hence, patch Euler class (1). We shall first present the idea in the simple setting of a Kagome model by considering the $(K,K^{\prime})$ nodes in between bands 1 and 2, where the $\mathcal{C}_{6}$ symmetry eliminates the complicating effects of dynamic phases, isolating the key physics at play. We will subsequently discuss more general settings, including the effects of DSs and dynamic phases.
Figure 2: a,b) Interferometry to distinguish relative frame charges of nodes in Kagome lattice. Up- and down-spin atoms are moved along paths I and II, respectively, and recombined. Opposite (a) [same (b)] frame charges result in a single spin [mixture of both spins] as in Table. 1. In (b2), interferometry paths enclose an additional node, crossing a DS (blue line). c) Illustration of the general interferometry. Geometric phases $\Delta\alpha_{n}-\Delta\alpha_{n+1}=0$ at the close of the loop indicates oppositely charged nodes, which will annihilate if brought together within the shaded area, i.e., vanishing patch Euler class.
@@
Assuming that the frame-rotation angle for the node I is $\phi$, the atoms following this path finish in the state $\big{(}c\mathinner{|{u_{1}}\rangle}-s\mathinner{|{u_{2}}\rangle}\big{)}\mathinner{|{\uparrow}\rangle}$ as in Eq. (Interferometry of non-Abelian band singularities and Euler class topology), where $c\equiv\cos{\phi}$ and $s\equiv\sin{\phi}$ and the momentum index is suppressed for simplicity. If the node on path II carries the same charge (in the chosen gauge), the frame features the same chirality on both nodes. However, since the path winds in the opposite direction ($-\theta$), the frame-rotation angle experienced on path II is $-\phi$. Conversely, if the nodes host opposite charges, then path II corresponds to a frame-rotation angle $\phi$  [^3] We consider these interferometry angles to be such that the magnitude of frame rotation angle is the same on both paths, which can be controlled e.g. using the techniques in Ref. [34] even for reduced lattice symmetries..
Therefore, considering the paths together, the final state becomes
@@
where the +()+(-)+ ( - ) for down-spins on path II corresponds to same (opposite) charges. Applying another $\pi/2$ pulse before closing the interferometry then yields $c\mathinner{|{u_{1}}\rangle}\mathinner{|{\uparrow}\rangle}-s\mathinner{|{u_{2}}\rangle}\mathinner{|{\downarrow}\rangle}$ for similarly-charged nodes, resulting in a mixture of both spin populations, $p_{\uparrow}=\cos^{2}{\left(\phi\right)}\mathrm{~{}and~{}}p_{\downarrow}=\sin^{2}{\left(\phi\right)}$. On the other hand, for oppositely-charged nodes, the final state $c\mathinner{|{u_{1}}\rangle}\mathinner{|{\uparrow}\rangle}-s\mathinner{|{u_{2}}\rangle}\mathinner{|{\uparrow}\rangle}$ consists only of up spins $p_{\uparrow}=1$. We demonstrate our interferometry technique along different paths in Fig. 2(a,b) with resulting populations given in Table 1, where the relative non-Abelian frame charges are indeed distinguished.
@@
(b1) $c\mathinner{|{u_{1}}\rangle}\mathinner{|{\uparrow}\rangle}-s\mathinner{|{u_{2}}\rangle}\mathinner{|{\downarrow}\rangle}$ $\cos^{2}{\theta/2}$ $\sin^{2}{\theta/2}$ same
(b2) $c\mathinner{|{u_{1}}\rangle}\mathinner{|{\downarrow}\rangle}-s\mathinner{|{u_{2}}\rangle}\mathinner{|{\uparrow}\rangle}$ $\sin^{2}{\theta/2}$ $\cos^{2}{\theta/2}$ same
Table 1: Interferometry results for paths depicted in Fig. 2(a,b). While opposite frame charges result in a single spin species, similar charges yield a mixture.
General configurations.— Our interferometry technique readily caters to more complex band structures. The dynamic phases along the two paths considered above (Fig. 2) are equal under $\mathcal{C}_{6}$ symmetry and cancel each other (see Supplemental Material (SM) [44]). In general when the nodes of interest are not positioned at high-symmetry points, dynamic phases must be accounted for. This can be achieved e.g. by measuring band energies along the interferometry with standard techniques such as band mapping [45, 31], while we present another method based on performing a reference interferometry loop with twice the acceleration in the SM [44].
Furthermore, there can be Dirac strings traversed by the interferometry loop, which change the sign of the relevant band eigenstates, leading to an additional $\pi$ geometric phase. While the DS crossed twice in Fig. 2a yields no net effect as in Table. 1  [^4] Equivalently, the Dirac string can be deformed past the node along path I and omitted, the single crossing on path I in Fig. 2b induces a $\pi$ phase shift between the spins and switches $p_{\uparrow}$ and $p_{\downarrow}$, but still entails a spin mixture for the similar-valued frame charges probed. We here give a general proof for generic DS configurations and show that our interferometry scheme determines whether the nodes can annihilate if brought together along the paths that close the interferometry loop (see Fig. 2c). Given the Euler class, this therefore reveals the relative frame charges and patch invariant (1) within this region.
We consider two Euler nodes between bands $n$ and $n+1$, and focus on the geometric phases $\alpha_{n(n+1)}^{\text{I/II}}$ in each band along the paths I and II, where the dynamic phases can be accounted for as before if required. We now analyze the effect of all possible DS configurations:
@@
3. When the wavepackets are brought together avoiding other nodes, there can be three types of DSs crossed; within the Euler subspace and in the adjacent gaps, as shown in Fig. 2c, which result in changes in $\Delta\alpha_{n(n+1)}$ summarized in Table 2. Most importantly, these DSs would also alter the relative sign of the targeted frame charges if they were to be brought together themselves along the closing paths. Indeed, we see that each time $\Delta\alpha_{n+1}-\Delta\alpha_{n}$ changes by $\pi$, the relative charge of the nodes changes.
4. The nodes’ obstruction to annihilate within a region of the Brillouin zone is captured by the patch Euler class (1). We therefore conclude that the patch Euler class in the area enclosing the second half of the interferometry [i.e. post-node branches] vanishes $\chi_{n,n+1}=0$ (is finite, $\chi_{n,n+1}=1$) if $|\Delta\alpha_{n+1}-\Delta\alpha_{n}|=0\,(\pi)$. Corresponding spin populations (after a $\pi/2$-pulse upon closing the interferometry [44]) are given in Table 3.
We here note that evaluating the Euler class is superfluous for two nodes separated by a reciprocal lattice vector. While we illustrate the relative charges using the simplest Kagome setting in Fig. 2b, our scheme applies to generic situations. We validate this by employing a Kagome model with next-nearest-neighbour tunneling terms [20] which host several more singularities of various charges, see SM [44]. Furthermore, under linear periodic driving, this model features an anomalous Euler phase [29] where the non-trivial patch invariant arises by virtue of a DS in the anomalous Floquet gap, which we confirm to be successfully captured by our interferometry [44].
@@
Table 2: Possible DS crossings (Fig. 2c) after the targeted nodes, and corresponding effects on the phase difference in each band and frame charges if the nodes were to be moved.
@@
00 $\pi$ $\cos^{2}{\phi}$ $\sin^{2}{\phi}$ same
$\pi$ 00 $\sin^{2}{\phi}$ $\cos^{2}{\phi}$ same
Table 3: Spin populations resulting from phase differences between the two paths in bands $n$ and $n+1$.

### v2

@@ Abstract
In systems with a real Bloch Hamiltonian band nodes can be characterised by a non-Abelian frame-rotation charge. The ability of these band nodes to annihilate pairwise is path dependent, since by braiding nodes in adjacent gaps the sign of their charges can be changed. Here, we theoretically construct and numerically confirm two concrete methods to experimentally probe these non-Abelian braiding processes and charges in ultracold atomic systems. We consider a coherent superposition of two bands that can be created by moving atoms through the band singularities at some angle in momentum space. Analyzing the dependency of excitations on the frame charges, we demonstrate an interferometry scheme passing through two band nodes, which reveals the relative frame charges and allows for measuring the multi-gap topological invariant. The second method relies on a single wavepacket probing two nodes sequentially, where the frame charges can be determined from the band populations. Our results present a feasible avenue for measuring non-Abelian charges of band nodes and the direct experimental verification of braiding procedures, which can be applied in a variety of settings including the recently discovered anomalous non-Abelian phases arising under periodic driving.
@@
Band nodes also play an important role in single-gap topological phases. Weyl nodes act as sources of Berry flux [30], while higher order degeneracies require stabilization of crystalline symmetries and are often associated with topological invariants. Individually, their properties can be probed in different experiments [31, 32, 33, 34, 35, 36]. For example, the $\pi$-Berry flux of a Dirac cone has been measured through atomic momentum-space interferometry in optical lattices [32], and the winding of isolated linear or quadratic band touchings has been observed via exciting atoms by moving them through nodes [34]. Upon being elevated to multi-gap topologies, the hallmark gained by band singularities is non-Abelian frame charges that can induce a nontrivial Euler class. Consequently, a fundamental question arises whether there exists observable signatures of these non-Abelian charges, their braiding and this invariant.
@@
Non-Abelian Frame Charges and Euler Class.— For a real Hamiltonian, the eigenstates $|u_{n}(\bm{k})\rangle$ at quasimomentum $\bm{k}$ span an orthonormal triad or, in general, vielbein (‘frame’): The space of spectralized Hamiltonians is given by $O(N)/O(1)^{N}=O(N)/\mathbb{Z}_{2}^{N}$ for $n\in N$ bands, where $N>2$. The quotient ensures that flipping the signs of the eigenvectors does not change the Hamiltonian. Band singularities obstruct a unique assignment of the eigenstates and correspond to a $\pi$-rotation between the eigenstates that host the node (see Fig. 1b)  [^1] Resulting in the $\pi$-Berry flux of a Dirac cone.. Namely, the frame rotation ($\phi^{\mathrm{frame}}$) accumulates a $\pi$ Berry phase upon circling around a node, where band singularities act as the analogues of $\pi$ disclination vortices in biaxial nematics [38, 39]. While within the two-band subspace (a single ‘gap’) the nodes act as $\pi$-rotations, we note that $2\pi$ rotations in fact correspond to $-1$ as $\pi_{1}(SO(N))=\mathbb{Z}_{2}$, accumulated frame charges anticommute with those in the adjacent gaps. Focusing on three bands, the frame charges associated to the nodes $\pi_{1}(SO(3)/D_{2})=Q$ take values in the quaternion group $Q=\{\pm 1,\pm i,\pm j,\pm k\}$ [38]. Consequently, braiding band nodes in momentum space (which can be achieved e.g. using stress/strain [17, 40], temperature effects [41] or periodic driving [29]), converts their charges and can ensure that a specific band subspace hosts similarly-valued charges. In two dimensions, the resulting obstruction to annihilate these charges (nodes) between states $n$ and $(n+1)$ is quantified by the Euler class [17, 42],
@@
evaluated over any patch $\mathcal{D}$ encapsulating these nodes, [Fig. 1b]. Here, we define the Euler form $\mathrm{Eu}=\langle\partial_{k_{1}}u_{n}(\boldsymbol{k})|\partial_{k_{2}}u_{n+1}(\boldsymbol{k})\rangle-\langle\partial_{k_{2}}u_{n}(\boldsymbol{k})|\partial_{k_{1}}u_{n+1}(\boldsymbol{k})\rangle$ and associated connection one-form, $\mathcal{A}=\langle u_{n}(\boldsymbol{k})|\boldsymbol{\nabla}u_{n+1}(\boldsymbol{k})\rangle$.
@@
The braiding and Euler class can effectively be captured by tracking Dirac strings (DSs) [29, 42]. These are gauge objects connecting pairs of nodes in each gap (see Fig. 1) and represent the line across which the sign of the eigenstates change due to the $\pi$-Berry phase induced by the node. Hence, crossing a DS residing in an adjacent gap changes the sign of the frame charge, effectively encoding braiding rules of the band node charges [42].
@@
Model.— As a concrete example, we employ a $\mathcal{C}_{2}\mathcal{T}$-symmetric Kagome lattice, while our results apply to any system admitting an Euler description. The momentum-space Hamiltonian is written as
@@
$H(\bm{k})=-2J\sum_{\beta\neq\beta^{\prime}}\cos{\left(\bm{k}\cdot\bm{d}_{\beta\beta^{\prime}}\right)}c^{\dagger}_{\beta}c^{\phantom{\dagger}}_{\beta^{\prime}}+\sum_{\beta}\Delta_{\beta}c^{\dagger}_{\beta}c^{\phantom{\dagger}}_{\beta},$ (2)
@@
with nearest-neighbour hopping amplitudes $J$ along the three directions $\bm{d}_{\beta\beta^{\prime}}$ connecting three sublattices $\beta\in(A,B,C)$ and the annihilation (creation) operator $c^{(\dagger)}_{\beta}$ [29]. For vanishing sublattice offsets $\Delta_{\beta}$, there are two linear band touching points which carry opposite frame charges (e.g. ±$\pm i$, depicted with empty/filled markers in Fig. 1b), hence a vanishing patch Euler class. The quadratic band touching point in gap 2 harbours $\chi=1$, corresponding to two same-valued frame charges which we can assign to be $+j$  [^2] The frame-charge assignment to each gap is a gauge choice, where only the relative sign in a gap is physically meaningful and the corresponding Euler class is gauge invariant.. While quaternion frame charges act as ±$\pm\pi$ Berry fluxes in a given gap, conventional interferometry methods are insensitive to their signs and, hence, cannot a priori reveal the relative frame charges or their change after braiding.
[…]
@@
Method 1: Interferometry.— We employ a coherent superposition in the bands forming a singularity (i.e. the Euler subspace) to deduce the frame rotation around it relative to other nodes in this gap, hence, their relative frame charges. Consider moving a wavepacket in momentum space in band 1 $(\mathinner{|{u_{1,\mathrm{in}}(\bm{k})}\rangle})$, which enters and exits a band node at some angle $\phi^{\mathrm{path}}$ as in Fig. 1[34]. Between the in-going and out-going paths the eigenstates frame is rotated by an angle $\phi^{\mathrm{frame}}$. Therefore, after passing through the band touching point, atoms are in a coherent superposition [42],
@@
$=\cos{\phi^{\mathrm{frame}}}\mathinner{|{u_{1,\mathrm{out}}}\rangle}-\sin{\phi^{\mathrm{frame}}}\mathinner{|{u_{2,\mathrm{out}}}\rangle},$ (3)
given by the overlap with the final eigenstates $\mathinner{|{u_{1,\mathrm{out}}}\rangle}$ and $\mathinner{|{u_{2,\mathrm{out}}}\rangle}$, captured by the frame rotation $\phi^{\mathrm{frame}}$. Note that we here assume an adiabatic motion with respect to the third band, which can be generally satisfied due to the energetic separation (Fig. 1a) and the orthonormality of the eigenstates. While the required path precision is within reach of state-of-the-art experiments [34], the desired superposition in the Euler subspace governed by $\phi^{\mathrm{frame}}$ can be still created even if atoms might just miss the node realistically as we show in Supplemental Material (SM) [42]. For linear nodes, $\phi^{\mathrm{frame}}$ winds always by $\pi$ as $\phi^{\mathrm{path}}$ winds by $2\pi$. To demonstrate our scheme, we consider uniform and symmetric Hamiltonians (e.g. Kagome) for a smooth winding $\phi^{\mathrm{frame}}=\pm\phi^{\mathrm{path}}/2$, while our methods readily apply to generic nodes  [^3] In general, $\phi^{\mathrm{frame}}$ can be a generic function of $\phi^{\mathrm{path}}$ but always winds by $\pi$ as the latter winds by $2\pi$. See Ref. [42] for more details and examples..
When the system is characterized by an Euler class, the relative phase of the component excited into the second band depends on the chirality of the frame winding, i.e. the full dreibein that is naturally encoded by the Hamiltonian evolution. By analyzing two such excitations, we devise an interferometry of the Euler nodes to extract the relative non-Abelian frame charges and, hence, patch Euler class (1). We shall first present the idea in the simple setting of a Kagome model by considering $(K,K^{\prime})$ nodes, where the $\mathcal{C}_{6}$ symmetry eliminates complicating effects of dynamic phases, isolating the key physics at play. We subsequently discuss more general settings, including effects of DSs and dynamic phases.
Figure 2: a,b) Interferometry distinguishing relative frame charges of nodes in Kagome lattice. Up- and down-spin atoms are moved along paths I and II, respectively, and recombined. Opposite (a) [same (b)] frame charges result in a single spin [mixture of both spins] as in Table. 1. In (b2), interferometry paths enclose an additional node, crossing a DS (blue line). c) Illustration of the general interferometry (triangles, circles and squares indicate nodes between different bands as labeled). Final geometric phases $\Delta\alpha_{n}\!-\!\Delta\alpha_{n+1}\!=\!0$ indicate oppositely charged nodes, which annihilate if brought together within the shaded area, i.e., vanishing patch Euler class.
@@
Assuming that the frame-rotation angle for the node I is $\phi^{\mathrm{frame}}$, the atoms following this path finish in the state $\big{(}c\mathinner{|{u_{1}}\rangle}-s\mathinner{|{u_{2}}\rangle}\big{)}\mathinner{|{\uparrow}\rangle}$ as in Eq. (Interferometry of non-Abelian band singularities and Euler class topology), where $c\equiv\cos{\phi^{\mathrm{frame}}}$ and $s\equiv\sin{\phi^{\mathrm{frame}}}$ and the momentum index is suppressed for simplicity. If the node on path II carries the same charge with no DS in between, the frame features the same chirality on both nodes. However, since the path winds in the opposite direction ($-\phi^{\mathrm{path}}$), the frame-rotation angle experienced on path II is $-\phi^{\mathrm{frame}}$. Conversely, if the nodes have opposite charges, path II corresponds to a frame-rotation angle $\phi^{\mathrm{frame}}$  [^4] We consider these interferometry angles to be such that the magnitude of frame rotation angle is the same on both paths, which can be controlled e.g. using the techniques in Ref. [34] even for reduced lattice symmetries [42]..
Considering the paths together, the final state becomes
@@
where +()+(-)+ ( - ) for down-spins on path II corresponds to same (opposite) charges. Applying another $\pi/2$ pulse before closing the interferometry then yields $c\mathinner{|{u_{1}}\rangle}\mathinner{|{\uparrow}\rangle}-s\mathinner{|{u_{2}}\rangle}\mathinner{|{\downarrow}\rangle}$ for similarly-charged nodes, resulting in a mixture of both spin populations, $p_{\uparrow}=\cos^{2}{\left(\phi^{\mathrm{frame}}\right)}\mathrm{~{}and~{}}p_{\downarrow}=\sin^{2}{\left(\phi^{\mathrm{frame}}\right)}$. However, for oppositely-charged nodes, the final state $c\mathinner{|{u_{1}}\rangle}\mathinner{|{\uparrow}\rangle}-s\mathinner{|{u_{2}}\rangle}\mathinner{|{\uparrow}\rangle}$ consists only of up spins $p_{\uparrow}=1$. We demonstrate our interferometry technique along different paths in Fig. 2(a,b) with resulting populations given in Table 1 which are also numerically confirmed, where the relative non-Abelian frame charges are distinguished.
@@
(b1) $c\mathinner{|{u_{1}}\rangle}\mathinner{|{\uparrow}\rangle}-s\mathinner{|{u_{2}}\rangle}\mathinner{|{\downarrow}\rangle}$ $\cos^{2}{\phi^{\mathrm{path}}/2}$ $\sin^{2}{\phi^{\mathrm{path}}/2}$ same
(b2) $c\mathinner{|{u_{1}}\rangle}\mathinner{|{\downarrow}\rangle}-s\mathinner{|{u_{2}}\rangle}\mathinner{|{\uparrow}\rangle}$ $\sin^{2}{\phi^{\mathrm{path}}/2}$ $\cos^{2}{\phi^{\mathrm{path}}/2}$ same
Table 1: Interferometry results for paths in Fig. 2(a,b). While opposite frame charges result in a single spin species, similar charges yield a mixture.
General configurations.— Our interferometry technique readily caters to more complex band structures. The dynamic phases along the two paths considered above (Fig. 2) are equal under $\mathcal{C}_{6}$ symmetry and cancel in Table 1 (see SM [42]). In general when targeted nodes are not at high-symmetry points, dynamic phases must be accounted for. This can be achieved e.g. by measuring band energies along the interferometry with standard techniques like band mapping [46, 31]. We present another method based on performing a reference interferometry loop with twice the acceleration in SM [42].
Furthermore, there can be DSs traversed by the interferometry loop, which change the sign of the relevant eigenstates, leading to an additional $\pi$ geometric phase. While the DS crossed twice in Fig. 2a yields no net effect as in Table. 1  [^5] Equivalently, the Dirac string can be deformed past the node along path I and omitted, the single crossing on path I in Fig. 2b induces a $\pi$ phase shift between the spins and switches $p_{\uparrow}$ and $p_{\downarrow}$, but still entails a spin mixture for the similar-valued frame charges. We here give a proof for generic DS configurations that our interferometry scheme determines whether the nodes can annihilate if brought together along the paths closing the interferometry loop (Fig. 2c). This reveals the relative frame charges and gauge-invariant Euler class (1) within this region.
We consider two Euler nodes between bands $n$ and $n+1$, and focus on geometric phases $\alpha_{n(n+1)}^{\text{I/II}}$ in each band along paths I and II, where dynamic phases can be accounted for as before if required. We now analyze the effect of all possible DSs:
@@
3. When the wavepackets are brought together avoiding other nodes, there can be three types of DSs crossed; within the Euler subspace and above/below gap, see Fig. 2c: Crossing e.g. the yellow DS changes $\mathinner{|{u_{n+1}}\rangle}\rightarrow-\mathinner{|{u_{n+1}}\rangle}$ on path I, which reflects on $\Delta\alpha_{n+1}$, as summarized in Table 2 for all cases. If the node itself is moved across the yellow DS, its frame charge flips. Indeed, we see that each time $\Delta\alpha_{n+1}-\Delta\alpha_{n}$ changes by $\pi$, the relative charge of the nodes changes.
4. The nodes’ obstruction to annihilate within a region of the Brillouin zone is captured by the patch Euler class (1). We therefore conclude that the patch Euler class in the area enclosing the second half of the interferometry [i.e. post-node branches] vanishes $\chi_{n,n+1}=0$ (is finite, $\chi_{n,n+1}=1$) if $|\Delta\alpha_{n+1}-\Delta\alpha_{n}|=0\,(\pi)$. Corresponding spin populations (after a $\pi/2$-pulse upon closing the interferometry [42]) are given in Table 3 which are numerically verified.
We note that evaluating the Euler class is superfluous for two nodes separated by a reciprocal lattice vector. While we illustrate the relative charges using the simplest Kagome setting [Fig. 2], our scheme applies generically. We validate this by employing a Kagome model with next-nearest-neighbour tunneling terms [20] which host several more singularities of various charges, see SM [42]. Furthermore, under linear periodic driving, this model features an anomalous Euler phase [29] where the non-trivial patch invariant arises by virtue of a DS in the anomalous Floquet gap, which we confirm to be captured by our interferometry [42].
@@
Table 2: $\pi$-phase changes acquired by crossing possible DSs in Fig. 2c (shaded area) and corresponding effects on relative frame charges as explained in the text.
@@
00 $\pi$ $\cos^{2}{\phi^{\mathrm{frame}}}$ $\sin^{2}{\phi^{\mathrm{frame}}}$ same
$\pi$ 00 $\sin^{2}{\phi^{\mathrm{frame}}}$ $\cos^{2}{\phi^{\mathrm{frame}}}$ same
Table 3: Effects of possible phase changes [cf. Table. 2] on spin populations in bands $n,\,n+1$ after the interferometry: Relative frame charge depends only on $|\Delta\alpha_{n+1}-\Delta\alpha_{n}|$).